The demand curve is QD=500-1/2P


Find the point at which point elasticity is equal to -1.


1
Expert's answer
2021-11-07T19:46:25-0500

"Ed = -b\u00d7P\/Qd."

"-1\/2\u00d7(1000 - 2Q)\/Q = -1,"

(1000 - 2Q)/Q = 2,

1000 - 2Q = 2Q (Q is not equal zero),

4Q = 1000,

Q = 250 units,

P = 1000 - 2×250 = 500.